OpenStudy (anonymous):

Billy was counting the number of wheels and bike seats at the playground. He saw tricycles and bicycles. He counted 79 wheels and 28 seats. how many bicycles did he see?

OpenStudy (anonymous):

Well, lets just say for example, 18 of these seats belong to a bicycle, so therefore 18*2=36 & 79-36=43 , which only leaves 43 wheels. But since 43/3=14.333, since its impossible to get a decimal of existent things its definitely not that, so just try with more numbers, if that makes sense lol